# Voyaging into Unbounded Dynamic Scenes from a Single View

### v1

@@ Abstract
This paper studies the problem of generating an unbounded dynamic scene from a single view, which has wide applications in augmented/virtual reality and robotics. Since the scene is changing over time, different generated views need to be consistent with the underlying 3D motions. While previous works learn such consistency by training from multiple views, the generated scene regions are bounded to be close to the training views with limited camera movements. To address this issue, we propose DynamicVoyager that reformulates the dynamic scene generation as a scene outpainting process for new dynamic content. As 2D outpainting models can hardly generate 3D consistent motions from only 2D pixels at a single view, we consider pixels as rays to enrich the pixel input with the ray context, so that the 3D motion consistency can be learned from the ray information. More specifically, we first map the single-view video input to a dynamic point cloud with the estimated video depths. Then we render the partial video at a novel view and outpaint the video with ray contexts from the point cloud to generate 3D consistent motions. We employ the outpainted video to update the point cloud, which is used for scene outpainting from future novel views. Experiments show that our model is able to generate unbounded scenes with consistent motions along fly-through cameras, and the generated contents can be controlled with scene prompts.
@@ 1 Introduction
Scene generation [57, 58, 31] aims to create a virtual scene with 3D representations such as point clouds [57, 17], meshes [8], neural radiance fields [32, 43] and 3D Gaussians [21]. It has many applications in augmented reality (AR), virtual reality (VR) [5, 4] and robotics [9]. While recent studies have achieved significant progress in generating a static scene [27, 58, 57] from a single image by leveraging the power of diffusion models [3, 41], these approaches cannot generate scenes with vivid dynamic contents, which fail to meet downstream needs. For example, AR/VR game designers need to build a dynamic scene for the players to explore and interact. Likewise, robotics researchers often require a virtual scene containing commonly seen dynamic objects in natural environments for self-exploration embodied agents [52, 11].
Generating dynamic scenes from a single view is a challenging task. The generated dynamic contents (e.g., hands waving, rivers flowing) need to maintain 3D consistency: when we observe the generated dynamic scenes from two different views, the observed motions should have 3D correspondence between those two views. As we can only observe 2D motions on the image plane from the 2D image pixels at one single view, there is a 2D-to-3D ambiguity when generating 3D consistent motions at a novel view from these pixels. Previous works [50, 42, 56, 26, 28] address this challenge by learning from multiple views surrounding the dynamic scenes, so that the 3D motion information can be implicitly learned from the 2D motions observed in these views.
However, while these scene generation models trained on surrounding views successfully generate consistent 3D motions, the generated dynamic regions are strictly bounded by the input views with a limited range of camera motions, as exemplified in Figure 2. Such limitation severely restrains the models from creating unbounded dynamic scenes, which have an increasing demand in AR/VR and robotics. This further limits the development of controllable scene generation with human guidance. A natural question is raised: can we generate an unbounded dynamic scene from a single view?
In this paper, we provide a positive answer to this question. We present DynamicVoyager, which reformulates dynamic scene generation as a scene outpainting process, enabling the creation of new dynamic content in the unseen areas. In this way, our generated scenes can be explored from a single view to any place with fly-through camera trajectories. Note that traditional 2D outpainting models can hardly generate 3D consistent object motions by only leveraging 2D image pixel input at a single view, as 2D pixels in consecutive frames contain limited 3D motion information. To address this, we consider pixels as rays to enrich the pixel input with the context of camera rays, so that the 3D motion information can be learned from the ray context information in the 3D space.
@@
More specifically, for an image input, we exploit image-to-video diffusion models [54, 16] to synthesize a fixed pose video from that image. Then for the video input, we first estimate the video depth maps and backproject the fixed viewpoint video frames into 3D space as dynamic point clouds with the estimated depths. After that, we render a partial video from the reconstructed point clouds at a novel view. By viewing pixels as rays, we sample ray depth maps of overlapping regions between two views, and compute the distance between rays and the point clouds as complementary 3D information for the unseen areas. This allows us to outpaint the unseen regions in the partial video with the sampled ray depths and ray distance maps to generate 3D consistent motions. Finally, we estimate the depth maps of the outpainted video and update the unseen regions of the dynamic point clouds.Figure 1 visualizes examples of dynamic scenes generated by our method using input from a video with a fixed camera pose or a single image. Experiments demonstrate that our model effectively generates unbounded dynamic scenes with 3D consistent motions along fly-through camera trajectories and controls the outpainting with scene prompts.
@@ 2 Related Work
Static scene generation. With the development of vision foundation models [22, 47, 37, 20, 40], many works have started to generate static scenes from input images [10, 2, 7] or text prompts [31, 15]. Early efforts focused on indoor scene generation [15, 7, 2, 25] or object-centric scene generation [10]. These methods firstly generate multiview images of the target scene and fuse the multiview images by using 3DGS [21] or NeRF [32, 43, 44] representations. However, they can only generate static scenes with limited ranges of camera motions. When the camera moves drastically, these methods fail to generate unseen regions with consistent and clear scene structures. To overcome this issue, later studies have started to explore perpetual view generation [30] that allows larger camera movements [35, 58, 57, 30, 27, 19]. Notably, WonderJourney [58] and WonderWorld [57] propose to generate an infinite world with diverse contents by using a combination of state-of-the-art depth estimation [37], image segmentation [22] and object detection models [20]. However, these methods can only generate diverse scenes with static objects. In this paper, we target the problem of generating dynamic scenes with new dynamic contents with fly-through camera trajectories.
@@ 3 Approach
In this section, we propose DynamicVoyager for generating unbounded dynamic scenes, which includes three steps as summarized in Figure 3 and below.
  • 3.1 Dynamic Scene Initialization) DynamicVoyager accepts either a video with a fixed camera pose or an image with a prompt describing object motions; in the latter case, it first generates a video with a fixed pose. A depth map is then estimated from each frame of the video. We segment the video into a dynamic foreground and a static background, reconstruct the point clouds separately, and aggregate these into one dynamic point cloud.
  • 3.2 Ray Outpainting) At a new given pose, the goal is to generate a video consistent with the existing dynamic point cloud and a given scene prompt. We first rasterize the point cloud at the new pose, which is consistent by construction. That said, the rasterization only fills in part of the image, so the question is how to outpaint the rest consistently. To do so, we propose to guide the outpainting process with 3D information by providing the model with ray depth in the observed region and the distance of rays to the point cloud in the unseen region.
  • 3.3 Dynamic Scene Update) We update the dynamic point cloud via the video from the new pose. We repeat §3.2 and §3.3 to iteratively update the dynamic point cloud from novel poses.
@@ 3.1 Dynamic Scene Initialization
We begin by describing how to generate an initial dynamic point cloud given a video at a fixed camera pose $\bm{\Pi}^{(0)}$ or an image with a motion prompt. Notation-wise, a superscript $(i)$ denotes objects related to the $i^{\mathrm{th}}$ camera pose $\bm{\Pi}^{(i)}$.
@@
Given the generated or user-input fixed viewpoint video ${\bm{I}}^{(0)}$, we reconstruct the underlying dynamic scene, represented as 4D point cloud $\mathcal{P}^{(0)}=\{{\bm{p}}=({\bm{x}},t,{\bm{c}})\}$, where ${\bm{x}}\in\mathbb{R}^{3}$, $t\in\mathbb{R}$ and ${\bm{c}}\in\mathbb{R}^{3}$ denote the 3D position, timestamp and color respectively. Since the objects in the video foreground and the scene in the background have different dynamic natures, our framework reconstructs point cloud $\mathcal{P}^{\mathrm{(0,f)}}$ for the foreground scene and $\mathcal{P}^{\mathrm{(0,b)}}$ for background separately. To this end, we extract the binary video foreground masks ${\bm{M}}^{(0)}=\{{\bm{M}}^{(0)}_{t}\in\{0,1\}^{h\times w}\}_{t=0}^{N-1}$ with the foreground segmentation method [20] to segment the video ${\bm{I}}^{(0)}$ into foreground video ${\bm{I}}^{(0,\mathrm{f})}=\{{\bm{I}}^{(0,\mathrm{f})}_{t}\}_{t=0}^{N-1}$ and background video ${\bm{I}}^{(0,\mathrm{b})}=\{{\bm{I}}^{(0,\mathrm{b})}_{t}\}_{t=0}^{N-1}$ We employ the depth model [37] on video ${\bm{I}}^{\mathrm{(0)}}$ to obtain the video depth maps ${\bm{D}}^{(0)}=\{{\bm{D}}^{(0)}_{t}\in\mathbb{R}^{h\times w}\}_{t=0}^{N-1}$. Next, we describe the foreground and background scene initialization steps.
@@
Scene background initialization. Since the backgrounds do not have motions in the dynamic scene, we hypothesize that the pixel at the same location on the different video frames has the same 3D position at different timestamps. More specifically, let ${\bm{D}}^{(0)}_{t}(x,y),{\bm{M}}^{(0)}_{t}(x,y)$ denote the depth and foreground mask values at $x^{\mathrm{th}}$ row and $y^{\mathrm{th}}$ column of the video frame ${\bm{I}}^{(0)}_{t}$. The background depth can be element-wise calculated by the following equation,
@@ 3.2 Ray Outpainting at a Novel View
We now need to outpaint the scene at a given novel pose $\bm{\Pi}^{(1)}$. We follow the assumption of previous works [57, 58, 23] that the view of $\bm{\Pi}^{(1)}$ overlaps moderately with that of $\bm{\Pi}^{(0)}$ to allow for outpainting a substantial portion of the scene. The question is, how can we outpaint a dynamic scene containing consistent motions with the existing dynamic point cloud $\mathcal{P}^{(0)}$?
We first rasterize the point cloud at $\bm{\Pi}^{(1)}$, so that the observed region in the resulting partial video has obviously consistent motions with the point cloud. To handle the unseen region, a naive proposal is to directly outpaint the partial video. However, this typically leads to inconsistencies across the boundary of observed and unseen regions, as we visualize in Figure 5. Our intuition is that pixels in each video frame only provide 2D information which is not enough to reconstruct the 3D motion. Therefore, we consider pixels as rays: during the rasterization, we save the ray depth map aligned with each pixel to provide 3D ray information in the observed region; in the unseen region, we compute the distance between the rays (from the camera origin through the pixels) and the point cloud, to serve as guidance on how much the generated pixel should respect the existing point cloud.
Ray information sampling at the observed regions. Given a novel pose $\bm{\Pi}^{(1)}$, we first rasterize the dynamic point cloud $\mathcal{P}^{(0)}$: for every time $t\in\{0,\cdots,N-1\}$, we do
@@
Here, $\varphi$ denotes the image rasterization function. ${\bm{\hat{M}}}^{(1)}_{t}$ is a binary mask where ${\bm{\hat{M}}}^{(1)}_{t}(x,y)=1$ if at least one 3D point from $\mathcal{P}^{(0)}_{t}$ hits the image pixel at the $x^{\mathrm{th}}$ row and $y^{\mathrm{th}}$ column during rasterization and 00 otherwise. Correspondingly, ${\bm{\hat{I}}}^{(1)}_{t}$ is the rasterized partial video and ${\bm{\hat{D}}}^{(1)}_{t}$ the rasterized ray depth map (to be distinguished from the depth map generated by depth models as in §3.1). A hat on the notation emphasizes that the signals are not observed for all locations, and we denote ${\bm{\hat{D}}}^{(1)}=\{{\bm{\hat{D}}}^{(1)}_{t}\}_{t=0}^{N-1}$, ${\bm{\hat{M}}}^{(1)}=\{{\bm{\hat{M}}}^{(1)}_{t}\}_{t=0}^{N-1}$ and ${\bm{\hat{I}}}^{(1)}=\{{\bm{\hat{I}}}^{(1)}_{t}\}_{t=0}^{N-1}$. A few remarks are in order. First, by doing the rasterization, the observed regions are obviously consistent with the point cloud. Moreover, the ray depth maps encode the 3D information of the observed scene, which as we shall see will be included as guidance for outpainting. For the regions not observed at $\bm{\Pi}^{(1)}$, we introduce the distance between rays and the point cloud in the following to complement the 3D information in these regions.
@@
Figure 6: Voyaging into unbounded dynamic scenes with long camera trajectories from a single image or fixed viewpoint video. We iteratively show the unseen regions through large camera motions, the dynamic scene outpainting results and the consistent motions from novel views. By enriching the pixel information with ray contexts, our method successfully builds large dynamic scenes from input views.
@@
where ${\bm{r}}^{(1)}(x,y)\in\mathbb{R}^{3}$ is the unit-norm ray vector starting from the origin of $\bm{\Pi}^{(1)}$ pointing at pixel $(x,y)$, and ${\bm{o}}^{(1)}$ is the camera center of $\bm{\Pi}^{(1)}$, both in the reference coordinate system of $\bm{\Pi}^{(0)}$. As we will see, such a distance will be used to guide the outpainting.
Ray outpainting. To outpaint the dynamic scene at the novel pose $\bm{\Pi}^{(1)}$ from the partial video, we desire a few properties for the outpainted video: 1) the outpainted video needs to have a static pose fixed at $\bm{\Pi}^{(1)}$ and be consistent with the existing dynamic point cloud in the observed region, 2) the user should be able to control the generation in the outpainted region (i.e., the unseen region), and 3) the motions in the outpainted region need to be consistent with the motions in the observed region. To fulfill properties 1) and 2), we design our video outpainting model to take as input the extracted partial video ${\bm{\hat{I}}}^{(1)}_{t}$, as well as a prompt narrating the desired motions in the generated scene. For property 3), we propose to employ ray depth map ${\bm{\hat{D}}}^{(1)}$ and ray distance map ${\bm{\bar{D}}}^{(1)}$ as guidance on 3D information in the observed region and how much the outpainted ray in the unseen region should respect the existing point cloud. To sum up, our video outpainting model accepts the partial video ${\bm{\hat{I}}}^{(1)}$, scene prompt $\mathbf{s}$, ray depth map ${\bm{\hat{D}}}^{(1)}$, ray distance map ${\bm{\bar{D}}}^{(1)}$ and generates an outpainted video ${\bm{I}}^{(1)}=\{{\bm{I}}^{(1)}_{t}\}_{t=0}^{N-1}$. Next we describe how to train such an video outpainting model.
@@
Training outpainting model. As we need the ray information to train our video outpainting model, we extract depth maps of the training videos and initialize a camera with the same intrinsics of $\bm{\Pi}^{(0)}$ to calculate the ray information. More specifically, given a training video ${\bm{I}}$, we extract the video depth maps ${\bm{D}}$ with the depth model [37]. We apply a random video mask to the video and the depth maps to generate partial video ${\bm{\hat{I}}}$ and partial depth maps ${\bm{\hat{D}}}$. Then we initialize a fixed camera pose $\bm{\Pi}$ at the world center whose intrinsics are the same as $\bm{\Pi}^{(0)}$ and map the partial video to the point cloud with the partial depths. We employ (7) to calculate the distance ${\bm{\bar{D}}}$ between a ray and the point cloud. We progressively add noise to the video ${\bm{\hat{I}}}$ to generate noisy video $z_{\tau}$ where $\tau$ is the noise step. We employ ControlNet [59] to introduce the ray information into the video outpainting model. Let $\theta_{\mathrm{v}},\theta_{\mathrm{c}}$ denote the training parameters of the video diffusion model and ControlNet model. The training objective is
$\mathcal{L}=\mathbb{E}_{z_{0},{\bm{\hat{I}}},\tau,{\bm{\hat{D}}},{\bm{\bar{D}}},\mathbf{s},\epsilon}||\epsilon-\epsilon_{\theta_{\mathrm{v}},\theta_{\mathrm{c}}}({\bm{\hat{I}}},z_{\tau},\tau,{\bm{\hat{D}}},{\bm{\bar{D}}},\mathbf{s})||^{2}_{2},$ (8)
where $\epsilon\sim\mathcal{N}(0,1)$ is the diffusion noise. We train the outpainting model in an end-to-end manner.

### v2

# Voyaging into Perpetual Dynamic Scenes from a Single View

@@ Abstract
The problem of generating a perpetual dynamic scene from a single view is an important problem with widespread applications in augmented and virtual reality, and robotics. However, since dynamic scenes regularly change over time, a key challenge is to ensure that different generated views be consistent with the underlying 3D motions. Prior work learns such consistency by training on multiple views, but the generated scene regions often interpolate between training views and fail to generate perpetual views. To address this issue, we propose DynamicVoyager, which reformulates dynamic scene generation as a scene outpainting problem with new dynamic content. As 2D outpainting models struggle at generating 3D consistent motions from a single 2D view, we enrich 2D pixels with information from their 3D rays that facilitates learning of 3D motion consistency. More specifically, we first map the single-view video input to a dynamic point cloud using the estimated video depths. We then render a partial video of the point cloud from a novel view and outpaint the missing regions using ray information (e.g., the distance from a ray to the point cloud) to generate 3D consistent motions. Next, we use the outpainted video to update the point cloud, which is used for outpainting the scene from future novel views. Moreover, we can control the generated content with the input text prompt. Experiments show that our model can generate perpetual scenes with consistent motions along fly-through cameras. Project page: https://tianfr.github.io/DynamicVoyager.
@@ 1 Introduction
Perpetual scene generation [30, 27, 58] aims to create a virtual 3D scene as the camera moves along arbitrary trajectories, typically starting from a single view observation. While recent studies [27, 58, 57] have achieved significant progress in perpetual generation of static scenes by leveraging image outpainting models [41, 24, 23], these approaches cannot generate scenes with dynamic content (e.g., waving hands, flowing rivers). Such dynamic content has important applications in augmented reality (AR), virtual reality (VR) [4, 3] and robotics [7]. For example, AR/VR game designers need to build a perpetual dynamic scene for players to explore and interact, while roboticists use virtual scenes containing commonly seen dynamic objects in natural environments for training embodied agents [52, 9] by self-exploration.
A key challenge in generating perpetual dynamic scenes is to ensure that the generated dynamic content has 3D consistent motions: i.e., the motions observed in any two views must correspond to the same underlying 3D dynamics. Generating consistent motions from a single view is inherently ambiguous due to the limited 3D motion information contained in a 2D image. Previous dynamic scene generation methods [50, 42, 56, 26, 28] address this challenge by learning from multiple views surrounding the scenes so that 3D motion information can be implicitly learned from the 2D motions observed in these views. However, such a learning strategy severely restrains models from creating perpetual dynamic content, because the generated dynamic regions often interpolate between training views and fail to generate perpetual views, as exemplified in Figure 2.
In this paper, we propose DynamicVoyager, a novel approach to perpetual dynamic scene generation that reformulates the task as a scene outpainting problem, enabling the synthesis of new dynamic content in previously unseen regions of the scene. Rather than relying on multiple 2D views as input, DynamicVoyager learns 3D-consistent motion dynamics from single-view observations by treating pixels as rays, thereby enriching each pixel with the contextual information of its corresponding camera ray.
More specifically, given an input image, we exploit image-to-video diffusion models [54, 14] to synthesize a fixed pose video from that image. To initialize a dynamic scene from the video input, we estimate the video depth maps and backproject the fixed viewpoint video frames into 3D space as dynamic point clouds with the estimated depths. After that, we move the camera to a partially unseen area of the scene and rasterize the reconstructed point clouds into an incomplete video from that view. To outpaint the video with 3D consistent motions, we consider pixels as rays to complement the video input with 3D scene information. For pixels in the visible area, we obtain 3D motion information by sampling ray depth maps from the point clouds. For pixels in the unseen area, we backproject rays from pixels and use the distance between these rays and the point cloud to infer the 3D spatial relationship between the unseen area and the visible area. We then outpaint the incomplete video with the sampled ray depth and ray distance maps for the visible and unseen parts, respectively. Finally, we estimate the depth maps of the outpainted video and update the unseen area of the dynamic point clouds. Figure 1 visualizes examples of dynamic scenes generated by our method using input from a video with a fixed camera pose or a single image. It also shows that we can control the perpetual generation with input scene prompts.
@@
To summarize, our contributions are as follows:
  1. 1.
    We reformulate the dynamic scene generation problem as a scene outpainting problem, so that our generated scene can be explored from a single view to any place through fly-through camera trajectories.
  2. 2.
    We generate view-consistent motions in 3D space by treating pixels as rays, which allows us to use 3D information to enrich the outpainting model.
  3. 3.
    We present experiments showing that, unlike other bounded scene generation methods, our model can generate perpetual dynamic scenes and control the generation with scene prompts.
@@ 2 Related Work
Static scene generation. With the development of vision foundation models [20, 47, 37, 18, 40], many works have started to generate static scenes from input images [8, 2, 6] or text prompts [31, 13]. Early efforts focused on indoor scene generation [13, 6, 2, 25] or object-centric scene generation [8]. These methods firstly generate multiview images of the target scene and fuse the multiview images by using 3DGS [19] or NeRF [32, 43, 44] representations. However, they can only generate static scenes with limited ranges of camera motions. When the camera moves drastically, these methods fail to generate unseen regions with consistent and clear scene structures. To overcome this issue, later studies have started to explore perpetual scene generation [30] that allows larger camera movements [35, 58, 57, 30, 27, 17]. Notably, WonderJourney [58] and WonderWorld [57] propose to generate an infinite world with diverse contents by using a combination of state-of-the-art depth estimation [37], image segmentation [20] and object detection models [18]. However, these methods can only generate diverse scenes with static objects. In this paper, we target the problem of generating dynamic scenes with new dynamic contents with fly-through camera trajectories.
@@ 3 Approach
In this section, we present the proposed DynamicVoyager approach for generating perpetual dynamic scenes, which is summarized in Figure 3. DynamicVoyager consists of three steps, described in §3.1, §3.2 and §3.3, respectively:
  • Dynamic Scene Initialization: DynamicVoyager accepts either a video with a fixed camera pose or an image with a prompt describing object motions as input—in the latter case, it first generates a video with a fixed pose. A depth map is then estimated from each frame of the video. We segment the video into a dynamic foreground and a static background, reconstruct the point clouds separately, and aggregate these into one dynamic point cloud.
  • Ray Outpainting: At a new given pose, the goal is to generate a video consistent with the existing dynamic point cloud and a given scene prompt. We first rasterize the point cloud at the new pose, which is consistent by construction. That said, the rasterization only fills in part of the image, so the question is how to outpaint the rest consistently. To do so, we propose to guide the outpainting process with 3D information by providing the model with ray depth in the observed region and the distance of rays to the point cloud in the unseen region.
  • Dynamic Scene Update: We update the dynamic point cloud using the depth maps of the video frames from the new pose. We repeat the steps in §3.2 and §3.3 to iteratively update the dynamic point cloud from novel poses.
@@ 3.1 Dynamic Scene Initialization
[…]
@@
Given the generated or user-input fixed viewpoint video ${\bm{I}}^{(0)}$, we reconstruct the underlying dynamic scene, represented as a 4D point cloud $\mathcal{P}^{(0)}=\{{\bm{p}}=({\bm{x}},t,{\bm{c}})\}$, where ${\bm{x}}\in\mathbb{R}^{3}$, $t\in\mathbb{R}$ and ${\bm{c}}\in\mathbb{R}^{3}$ denote the 3D position, timestamp and color respectively. Since the objects in the video foreground and the scene in the background have different dynamic natures, our framework reconstructs point cloud $\mathcal{P}^{\mathrm{(0,f)}}$ for the foreground scene and $\mathcal{P}^{\mathrm{(0,b)}}$ for background separately. To this end, we extract the binary video foreground masks ${\bm{M}}^{(0)}=\{{\bm{M}}^{(0)}_{t}\in\{0,1\}^{h\times w}\}_{t=0}^{N-1}$ with the foreground segmentation method [18] to segment the video ${\bm{I}}^{(0)}$ into foreground video ${\bm{I}}^{(0,\mathrm{f})}=\{{\bm{I}}^{(0,\mathrm{f})}_{t}\}_{t=0}^{N-1}$ and background video ${\bm{I}}^{(0,\mathrm{b})}=\{{\bm{I}}^{(0,\mathrm{b})}_{t}\}_{t=0}^{N-1}$ We employ the depth model [37] on video ${\bm{I}}^{\mathrm{(0)}}$ to obtain the video depth maps ${\bm{D}}^{(0)}=\{{\bm{D}}^{(0)}_{t}\in\mathbb{R}^{h\times w}\}_{t=0}^{N-1}$. Next, we describe the foreground and background scene initialization steps.
@@
[…]
@@ 3.2 Ray Outpainting at a Novel View
To outpaint the scene at a given novel pose $\bm{\Pi}^{(1)}$, we follow the assumption of previous works [57, 58, 21] that the view of $\bm{\Pi}^{(1)}$ overlaps moderately with that of $\bm{\Pi}^{(0)}$ to allow for outpainting a substantial portion of the scene. The question is, how can we outpaint a dynamic scene containing consistent motions with the existing dynamic point cloud $\mathcal{P}^{(0)}$?
To address this question, we first rasterize the point cloud at $\bm{\Pi}^{(1)}$. As shown in the top row of Figure 5, we obtain a partial video whose observed region is the projection of the portion of the point cloud visible from $\bm{\Pi}^{(1)}$. To handle the unseen region, a naive approach would be to directly outpaint the partial video. However, this typically leads to inconsistencies across the boundary of observed and unseen regions, as shown in Figure 5. Our intuition is that pixels in each video frame only provide 2D information, which is not enough to reconstruct the 3D motion. To address this issue, we consider the 3D ray from the camera origin through each pixel. If the pixel is in the observed region, we compute its depth. If the pixel is in the unseen region, we compute the distance from its ray to the point cloud. Both the depth and distance maps serve as guidance for how to outpaint a pixel while respecting information from the existing point cloud.
Ray information sampling in observed region. Given a novel pose $\bm{\Pi}^{(1)}$, we first rasterize the dynamic point cloud $\mathcal{P}^{(0)}$: at every time step $t\in\{0,\cdots,N-1\}$, we compute
@@
[…]
@@
Figure 6: Voyaging into perpetual dynamic scenes with long camera trajectories from a single image or fixed viewpoint video. We iteratively show the unseen regions through large camera motions, the dynamic scene outpainting results and the consistent motions from novel views. By enriching the pixel information with ray contexts, our method successfully builds large dynamic scenes from input views.
@@
[…]
Ray outpainting. To outpaint the dynamic scene at the novel pose $\bm{\Pi}^{(1)}$ from the partial video, we desire a few properties for the outpainted video: 1) the outpainted video needs to have a static pose fixed at $\bm{\Pi}^{(1)}$ and be consistent with the existing dynamic point cloud in the observed region, 2) the user should be able to control the generation in the outpainted region (i.e., the unseen region), and 3) the motions in the outpainted region need to be consistent with the motions in the observed region. To fulfill properties 1) and 2), we design our video outpainting model to take as input the extracted partial video ${\bm{\hat{I}}}^{(1)}_{t}$, as well as a prompt narrating the desired motions in the generated scene. For property 3), we propose to employ ray depth map ${\bm{\hat{D}}}^{(1)}$ and ray distance map ${\bm{\bar{D}}}^{(1)}$ as guidance on 3D information in the observed region and how much the outpainted ray in the unseen region should respect the existing point cloud. To sum up, our video outpainting model accepts the partial video ${\bm{\hat{I}}}^{(1)}$, scene prompt $\mathbf{s}$, ray depth map ${\bm{\hat{D}}}^{(1)}$, ray distance map ${\bm{\bar{D}}}^{(1)}$ and generates an outpainted video ${\bm{I}}^{(1)}=\{{\bm{I}}^{(1)}_{t}\}_{t=0}^{N-1}$. Next we describe how to train such a model.
@@
Training the outpainting model. We collect a small video dataset for training (details in supplementary material), and reconstruct point clouds from it to compute ray information for supervision. More specifically, given a training video ${\bm{I}}$, we extract depth maps ${\bm{D}}$ using the depth model [37] and backproject the video from a camera $\bm{\Pi}$ at the world center to reconstruct a point cloud. We then move the camera closer to the scene and filter out the 3D points that are no longer visible from the current viewpoint. Then we move the camera back to the original camera pose $\bm{\Pi}$ and compute a partial video $^{\bm{\hat{I}}}$, a partial depth map $^{\bm{\hat{D}}}$, and a ray distance map ${\bm{\bar{D}}}$ following §3.2. To train the model, we add noise to ${\bm{I}}$ to obtain $z_{\tau}$ at step $\tau$, and employ ControlNet [59] to inject ray information. Let $\theta_{\mathrm{v}},\theta_{\mathrm{c}}$ denote the training parameters of the video diffusion model and ControlNet model. The training objective is
$\small\mathcal{L}=\mathbb{E}_{z_{0},{\bm{I}},{\bm{\hat{I}}},\tau,{\bm{\hat{D}}},{\bm{\bar{D}}},\mathbf{s},\epsilon}\big{\|}\epsilon-\epsilon_{\theta_{\mathrm{v}},\theta_{\mathrm{c}}}({\bm{I}},{\bm{\hat{I}}},z_{\tau},\tau,{\bm{\hat{D}}},{\bm{\bar{D}}},\mathbf{s})\big{\|}^{2}_{2},$ (8)
where $\epsilon\sim\mathcal{N}(0,1)$ is the diffusion noise.